The base of a triangle is two less than one-third of the height of a triangle. Find the area of the triangle as a function of the base.
Answer by flame8855(424)

let b be the base
let h be the height
b= h/3 -2 , 3b = h -6 , h=3b+6
Area= 1/2 (b)(h)=1/2 (b) (3b+6)=(3/2)b^2+3b
Area =(3/2)b^2+3b
yes send me ur question and I will help you sure :)